OpenStudy (anonymous):

prove this identity: (sinx+tanx)/(cosx+1)=tanx

hartnn (hartnn):

write tan x as sin x/ cos x

hartnn (hartnn):

what u get ?

OpenStudy (anonymous):

(sinx+(sinx/cosx))/(cosx+1) idk what to do after it

hartnn (hartnn):

lets work on numerator factor out sin x , what u get ?

OpenStudy (anonymous):

i did: (sinx+cosxsinx)/(cosx+1)

hartnn (hartnn):

u missed a cos x

hartnn (hartnn):

\(\large sin x+sinx/cos x= (sinxcosx+sinx)/cosx \\ \large=sinx(cos x+1)/cos x\) did u get this ?

hartnn (hartnn):

that was numerator only, the denominator still is 1+cos x

OpenStudy (anonymous):

howd you get: sinx(xosx+1)/cosx?

OpenStudy (anonymous):

cosx*

hartnn (hartnn):

i factored out sin x...

hartnn (hartnn):

\(a+b/c= (ac+b)/c\)